11. Alabama Power has preferred stock that pays an annual dividend of $9.44.lf the security has no maturity, what is its value to an investor who wishes to obtain a 10 percent rate of return? 0 a. $84.96 b. $104.89 c. $95.34 d. $94.40
Answer #1

Here the rate of return = 10% = 0.10 & the annual dividend = 9.44. Plug those into the PRESENT VALUE INTEREST FACTOR FOR ANNUITIES equation, which is:

PV = C * (1-((1/ (1+R))^n))/R . Here n is the maturity which is infinity as the security has no maturity, it is expect to pay dividend lifelong.

PV = 9.44 * (1-(1/(1.10))^(inf.))/.10

Notice 1/1.10 ~ .9090909 which is less than one. Multiplying a number less than 1 by itself an infinite amount of times will result in an infinitely small number, extremely close to zero. So

PV = 9.44 * (1-0)/.10

PV = 9.44/.10

PV = 94.4, option d is the correct answer.